OpenStudy (anonymous):

The volume of a rectangular prism is 288 ft3. If its height is 6 feet and its length is 8 feet, what is its width? A. 6 feet B. 8 feet C. 9 feet D. 12 feet

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (anonymous):

Not really... I'm fine with finding height but finding width is my weak spot.

OpenStudy (skullpatrol):

Recall the formula for the volume of a rectangular prism is V = l*w*h

OpenStudy (anonymous):

I am pretty sure it is 6...

OpenStudy (anonymous):

But try entering the possible answers into the equation.

OpenStudy (anonymous):

\[288=6*8*width\] \[288=48*width\] \[\frac{ 288 }{ 48 } = width\]

OpenStudy (anonymous):

Ok... Thanks. I'll try working off of the information everyone gave me.

OpenStudy (anonymous):

I believe the answer is 6...

OpenStudy (skullpatrol):

Correct :-)

OpenStudy (skullpatrol):

Don't forget your units.

OpenStudy (skullpatrol):

Not just 6. 6*(1 foot) = 6 feet. Unit^